OpenStudy (angelinaxoxo1):

Which proportion could be used to solve the problem below: 378 is 84% of what number? A. 84 over 100 = 378 over x B. 84 over 100 = x over 378 C. 378 over 100 = x over 378 D. 100 over 84 = 378 over x

OpenStudy (phi):

you should learn to translate the words into "math" 378 is 84% of what number? 378 is turns into = 84% turns into 84/100 or 0.84 of what number turns into " times x" put that all together: 378 = 84/100 * x

OpenStudy (phi):

that equation does not match any of your choices. so we have to fiddle with it divide both sides by x to get \[ \frac{378}{x}= \frac{84}{100} \]

OpenStudy (phi):

it still does not quite match, but you can swap sides in an equation: \[ \frac{84}{100}= \frac{378}{x} \]